Can someone please show me how to do this problem
Prove that the statement: 1/5+1/5^2+1/5^3+...+1/5^n=1/491-1/5^n) is true for all positive integers n. Show all your work
PLEASE HELP!!!!!
Thanks

Answer by oscargut(2103)

S=1/5+1/5^2+1/5^3+...+1/5^n
(1/5)S=1/5^2+1/5^3+...+1/5^n+1/5^(n+1)
S-(1/5)S = (4/5)S = 1/5 - (1/5)^(n+1)
S=1/4 - (5/4)(1/5)^(n+1)= 1/4(1-(1/5)^n)
i think you typed a 9 instead of a (